# COT2Align: Cross-Chain of Thought Distillation via Optimal Transport Alignment for Language Models with Different Tokenizers

### v1

@@ Abstract
COT2Align: Cross-Chain of Thought Distillation via Optimal Transport Alignment for Language Models with Different Tokenizers
Duc Anh Le1111Equal contribution, Tu Vu4111Equal contribution, Nam Le Hai1,† [^1] Equal contribution, Diep Thi-Ngoc Nguyen2,3, Linh Ngo Van1, Trung Le5, Thien Huu Nguyen6
1Hanoi University of Science and Technology, 2Oraichain Labs Inc., US, 4ByteDance Inc, 3VNU University of Engineering and Technology, 5Monash University, 6University of Oregon
@@ 4 Experiments
Methods KL ULD MinED DSKD GPT2-1.5B & Qwen1.5-1.8B $\rightarrow$ GPT2-120M Original $26.09$ $25.72$ $26.45$ $26.66$    + ours $\mathbf{31.36_{\pm 0.09}}$ $\mathbf{26.04}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.32}$ $\mathbf{26.81}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.36}$ $\mathbf{27.48}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.82}$ Llama2-7B & Mistral-7B $\rightarrow$ TinyLLaMA-1.1B Original $29.29$ $29.26$ $29.20$ $29.69$    + ours $\mathbf{29.99}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.7}$ $\mathbf{29.78}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.52}$ $\mathbf{30.16}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.96}$ $\mathbf{31.38}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 1.69}$ Qwen2.5-7B-Instruct $\rightarrow$ GPT2-1.5B Original - $27.65$ $28.71$ $28.90$    + ours - $\mathbf{28.33}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.68}$ $\mathbf{29.20}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.49}$ $\mathbf{30.77}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 1.87}$
@@ 4.3 Analysis
Table 3 provides an ablation study that highlights the individual contributions of each proposed component within COT2Align. Specifically, DSKD + COT2Align achieves the highest scores on all datasets, demonstrating its superiority in the knowledge distillation process (e.g., improvements on Dolly from 26.28 to 27.41 and on Dialogue Sum from 33.44 to 35.01). Both proposed losses individually enhance baseline performance (from 0.6 to over 1%), underscoring their effectiveness. Notably, $\mathcal{L}_{CRC}$ demonstrates greater improvement, highlighting the effectiveness of our constraint in guiding the CoT response to be more reliable. Moreover, even when applied solely to the last hidden layers, the improvement remains evident, further validating the effectiveness of our approach. Finally, the combination of these components in COT2Align yields best performance, highlighting their complementary nature and the ability to collectively optimize knowledge transfer. These results confirm the robustness of the proposed approach in enhancing the baseline KD method.
Methods Dolly Alpaca S-NI Dialogue Sum Mistral $\to$ TinyLLaMA DSKD 26.28$\pm$0.35 32.31$\pm$0.15 26.74$\pm$0.24 33.44$\pm$0.18    + $\mathcal{L}_{CST}$ 26.93$\pm$0.40 32.99$\pm$0.41 28.07$\pm$0.16 34.47$\pm$0.20    + $\mathcal{L}_{CRC}$ 26.93$\pm$0.46 33.18$\pm$0.58 29.51$\pm$0.24 34.39$\pm$0.24    + only hidden state layers 27.07$\pm$0.56 33.10$\pm$0.23 29.41$\pm$0.14 34.80$\pm$0.20 COT2Align 27.41$\pm$0.43 33.31$\pm$0.49 29.77$\pm$0.20 35.01$\pm$0.20

### v3

@@ 4 Experiments
Methods KL ULD MinED DSKD GPT2-1.5B & Qwen1.5-1.8B $\rightarrow$ GPT2-120M Original $26.09$ $25.72$ $26.45$ $26.66$    + ours $\mathbf{26.54}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.12}$ $\mathbf{26.04}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.32}$ $\mathbf{26.81}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.36}$ $\mathbf{27.48}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.82}$ Llama2-7B & Mistral-7B $\rightarrow$ TinyLLaMA-1.1B Original $29.29$ $29.26$ $29.20$ $29.69$    + ours $\mathbf{29.99}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.7}$ $\mathbf{29.78}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.52}$ $\mathbf{30.16}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.96}$ $\mathbf{31.38}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 1.69}$ Qwen2.5-7B-Instruct $\rightarrow$ GPT2-1.5B Original - $27.65$ $28.71$ $28.90$    + ours - $\mathbf{28.33}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.68}$ $\mathbf{29.20}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 0.49}$ $\mathbf{30.77}_{\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\uparrow 1.87}$
@@ 4.3 Analysis
Table 3 provides an ablation study that highlights the individual contributions of each proposed component within COT2Align. Specifically, DSKD + COT2Align achieves the highest scores on all datasets, demonstrating its superiority in the knowledge distillation process (e.g., improvements on Dolly from 26.28 to 27.41 and on Dialogue Sum from 33.44 to 35.01). Both proposed losses individually enhance baseline performance (from 0.6 to over 1%), underscoring their effectiveness. Notably, $\mathcal{L}_{CRC}$ demonstrates greater improvement, highlighting the effectiveness of our constraint in guiding the CoT response to be more reliable. We also conduct experiments using CoT augmentation into the original DSKD. The expermiental results are presented in the line "DSKD + only COT". It is evident that while CoT augmentation improves DSKD, its impact is not as significant as COT2Align.
Moreover, even when applied solely to the last hidden layers, the improvement remains evident, further validating the effectiveness of our approach. Finally, the combination of these components in COT2Align yields best performance, highlighting their complementary nature and the ability to collectively optimize knowledge transfer. These results confirm the robustness of the proposed approach in enhancing the baseline KD method.
Methods Dolly Alpaca S-NI Dialogue Sum Mistral $\to$ TinyLLaMA DSKD 26.28$\pm$0.35 32.31$\pm$0.15 26.74$\pm$0.24 33.44$\pm$0.18    + only COT 27.27$\pm$0.44 32.42$\pm$0.24 28.97$\pm$0.27 33.74$\pm$0.16    + $\mathcal{L}_{CST}$ 26.93$\pm$0.40 32.99$\pm$0.41 28.07$\pm$0.16 34.47$\pm$0.20    + $\mathcal{L}_{CRC}$ 26.93$\pm$0.46 33.18$\pm$0.58 29.51$\pm$0.24 34.39$\pm$0.24    + only hidden state layers 27.07$\pm$0.56 33.10$\pm$0.23 29.41$\pm$0.14 34.80$\pm$0.20 COT2Align 27.41$\pm$0.43 33.31$\pm$0.49 29.77$\pm$0.20 35.01$\pm$0.20